# Orbital orders under magnetic fields in cubic PrIr2Zn20

### v3

@@ Abstract
We study orbital orders in quadrupolar heavy-fermion compound PrIr2Zn20 under magnetic fields on the basis of the Landau theory. Assuming $E_{g}$ orbital (electric quadrupolar) orders in the cubic symmetry with the ordering wavenumber at the L points in the face-centered cubic lattice Brillouin zone as observed experimentally, we construct the Landau free energy and analyze it. We find that the unidentified high-temperature ordered phase under the magnetic field $\bm{H}\parallel[001]$ reported earlier can be interpreted as a consequence of the internal rotation in the orbital moments of f electrons at the Pr site. We also discuss the phase diagram for other magnetic-field directions and possible double-$\bm{q}$ quadrupolar orders in this system.
pacs:
Valid PACS appear here
@@ I Introduction
Strongly correlated electron systems with orbital moments have attracted significant attention in recent years. Tokura and Nagaosa (2000) Various multipolar orders appear in d- and f-electron-based materials and have been intensively studiedMydosh and Oppeneer (2014); Kuramoto et al. (2009); Chen et al. (2010). Among them, electric quadrupole orders are the most typical and extensively investigatedShiina et al. (1997); Onimaru et al. (2005); Chubukov et al. (2016). In addition, spin nematic orders have been also studied as well.Andreev and Grishchuk (1984); Läuchli et al. (2006); Tsunetsugu and Arikawa (2006); Jiang et al. (2023) The nature of orbital degrees of freedom leads to spatially anisotropic interactions, making the symmetry in orbital space generally anisotropic. Each type of orbital order has unique properties, and understanding these properties remains a key issue in condensed matter theory and experiment.
@@
Among the 1-2-20 compounds, PrIr2Zn20 is a prototypical material, exhibiting quadrupole ordering below 1 K and undergoing a phase transition to a superconducting state at 0.1 KOnimaru et al. (2011); Ishii et al. (2011); Onimaru et al. (2016). The early neutron scattering experimentIwasa et al. (2017) under an applied magnetic field $\bm{H}\parallel[\bar{1}10]$ suggested an ordering vector of $\bm{q}=(\pi,\pi,\pi)$, corresponding to the L point at the first Brillouin zone boundary. Here, we have set the lattice constant to unity. The analysis of induced magnetic moments indicates an orbital order of the $x^{2}-y^{2}$ type for $\bm{H}\parallel[\bar{1}10]$. Recent reinvestigations of the temperature–magnetic field phase diagram reveal strong anisotropy and highlight the existence of the high-temperature phaseKittaka et al. (2024) reported in earlier studiesOnimaru et al. (2011) across a wide range of magnetic field directions from [001] to [111], while it is absent for $\bm{H}\parallel[110]$.
In this paper, we analyze a phenomenological Landau model for PrIr2Zn20, incorporating $E_{g}$ quadrupole moments at the L points. We elucidate the nature of $E_{g}$ orbital ordering under magnetic fields and its dependence on field direction. We examine the order parameter suggested by the neutron scattering experiments for $\bm{H}\parallel[\bar{1}10]$, construct phase diagrams under applied magnetic fields, and compare them with experimental results. The ordering vectors at the L points reveal unique multiple-$\bm{q}$ orders, which differs from the previous studies about quadrupole orders at the X point ordering vectors in the face-center cubic lattice.Tsunetsugu et al. (2021); Hattori et al. (2023)
This paper is organized as follows. In Sec. II, we construct the Landau free energy for the $E_{g}$ quadrupole moments at wave vectors $\bm{q}=(\pi,\pi,\pi)$ and $(0,0,0)$. In Sec. III, we present the results of the Landau analysis for single-$q$ orbital orders, outlining the fundamental properties of the system. We then discuss multiple-$q$ orbital orders and the stability of various ordered phases in Sec. IV. In Sec. V, we compare our results with experimental data for PrIr2Zn20 and related materials. Finally, Sec. VI provides a summary of this paper.
@@ II Model
We start to introduce the exchange interactions between the $E_{g}$ orbital moments in Sec. II.1 and the Landau free energy for the order parameters at the L points in Sec. II.2. Mode coupling terms related to the order parameter at the L points and other degrees of freedom are discussed in Sec. II.5
Figure 1: Fourier transform of the exchange interaction for $J_{1}=1$ and (a) $(J_{2},J_{3},J_{4})=(0.0,0.0,0.0)$, (b) $(0.0,0.0,1.0)$, and (c) $(0.5,0.3,0.4)$. (d) The first Brillouin zone of the fcc lattice. The four L points are indicated by filled red circles and other symmetric points are indicated by filled black circles.
@@ II.1 Exchange interaction
Let us denote $E_{g}$ quadrupole moments at the position $\bm{r}=(x,y,z)$ in the diamond structure as $\bm{\phi}_{\bm{r}}^{\alpha}=(u_{\bm{r}}^{\alpha},v_{\bm{r}}^{\alpha})$, where $\alpha=A$ or $B$ is the sublattice index. We take one of the $A$ sublattice site at the origin $\bm{r}=(x,y,z)=(0,0,0)$ and its pair $B$ sublattice site in the unit cell at $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$, where we set the lattice constant $a=1$. $\bm{\phi}_{\bm{r}}^{\alpha}$ form a two-dimensional vector field as a classical variable. The two components are $u$ $\sim$ $2z^{2}$--$x^{2}$--$y^{2}$ and $v$ $\sim$ $\sqrt{3}$($x^{2}$--$y^{2}$). The Fourier transform in the wavenumber ${\bm{k}}$ space is defined as
${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}=\begin{bmatrix}u^{\alpha}_{{\bm{k}}}\\ v^{\alpha}_{{\bm{k}}}\end{bmatrix}=\frac{1}{\sqrt{N}}\sum_{\bm{r}}e^{-i{\bm{k}}\cdot\bm{r}}\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}.$ (1)
@@
$H=\sum_{{\bm{k}}}[u^{A}_{-{\bm{k}}},v^{A}_{-{\bm{k}}},u^{B}_{-{\bm{k}}},v^{B}_{-{\bm{k}}}]\begin{bmatrix}H^{AA}_{{\bm{k}}}&H^{AB}_{{\bm{k}}}\\ {H^{AB}_{{\bm{k}}}}^{\dagger}&H^{AA}_{{\bm{k}}}\end{bmatrix}\begin{bmatrix}u^{A}_{{\bm{k}}}\\ v^{A}_{{\bm{k}}}\\ u^{B}_{{\bm{k}}}\\ v^{B}_{{\bm{k}}}\end{bmatrix},$ (2)
where $H^{\alpha\beta}_{{\bm{k}}}$ is a $2\times 2$ matrix. We focus on the orders with the ordering wavenumber at the L points: ${\bm{k}}_{1}\equiv(\pi,\pi,\pi)$, ${\bm{k}}_{2}\equiv(-\pi,\pi,\pi)$, ${\bm{k}}_{3}\equiv(-\pi,-\pi,\pi)$, and ${\bm{k}}_{4}\equiv(\pi,-\pi,\pi)$. Since the exchange interactions at the L points are isotropic due to the symmetry, we retain such exchange interactions while neglect anisotropic ones. Although the anisotropic ones are generally present and can influence the ordering vectors and the excitation energy of quadrupole wave, it has no effect on the transition temperature for the order at the L points. Thus, a model without anisotropic interactions can be a good starting model as a minimal one. In this section, we consider exchange interactions up to the fourth neighbor interactions, while we will simplify the model in the later sections.
@@
$H^{AA}_{{\bm{k}}}$ $=[J_{2}\gamma_{2}({\bm{k}})+J_{4}\gamma_{4}({\bm{k}})]\hat{1},$ (3)
$H^{AB}_{{\bm{k}}}$ $=\frac{1}{2}[J_{1}\gamma_{1}({\bm{k}})+J_{3}\gamma_{3}({\bm{k}})]\hat{1}.$ (4)
@@
Here, $\hat{1}$ is a 2$\times$2 identity matrix. $J_{i}$’s are isotropic coupling constants. The expressions for $\gamma$’s are listed in Table 1. Note that $\gamma_{1}({\bm{k}})$ and $\gamma_{3}({\bm{k}})$ are complex. From the neutron experimental data, the ordering patterns are consistent with those expected for $J_{1}>0$ Iwasa et al. (2017).
@@
For describing the symmetry breaking at the L point modulation within our model, it is important to check which interactions are important. As will be explained later, four special wavenumbers at the $\Gamma$ point and the X points $\{(2\pi,0,0),(0,2\pi,0),(0,0,2\pi)\}$ are important other than the L point. First, among the same sub-lattice interactions, we note that the next-nearest interaction $\gamma_{2}$ vanishes at the four L points, while the fourth neighbor one $\gamma_{4}$ is finite as shown in Table 1. This means that positive $J_{4}$ favors the order at the L points while does not those at the $\Gamma$ and X points. For the inter sublattice ones, the nearest-neighbor interaction $\gamma_{1}$ and the third-neighbor one $\gamma_{3}$ are finite, while both vanish at the X points. In this sense, the effects of $J_{1}$ and $J_{3}$ are similar. Diagonalizing the matrix in Eq. (2), one obtains the eigenvalues $J({\bm{k}})$. Figure 1 shows $J({\bm{k}})$ for typical parameter sets with $J_{1}=1$ as a unit of energy. For the nearest-neighbor model with $J_{4}=0$ in Fig. 1(a), $J({\bm{k}})$ is smallest at the $\Gamma$ point, leading to a Néel type antifero quadrupole order. When $J_{4}$ is sufficiently large or $J_{2}$ is positive and not too large, the $J({\bm{k}})$ has minima at the L points as shown in Figs. 1(b) and 1(c), and these cases are situations we analyze in this paper.
@@
Table 1: exchange interaction coefficients. The values of $\gamma({\bm{k}})$’s at ${\bm{k}}=(0,0,0),(0,0,2\pi)$, ${\bm{k}}_{1}=(\pi,\pi,\pi)$, and the other L points are also shown. Here, we take the unit cell convention of $A$: (0,0,0) and $B$: $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$ and thus, the ${\bm{k}}_{1}$ is apparently inequivalent to the remaining ${\bm{k}}_{2,3,4}$. $\eta_{{\bm{k}}}\equiv e^{i(k_{x}+k_{y}+k_{z})/4}$, $\zeta^{\pm}_{{\bm{k}}}\equiv\pm 2(\cos\frac{k_{x}}{2}+\cos\frac{k_{y}}{2}+\cos\frac{k_{z}}{2})-3$. The mean-field second-order transition temperatures $T_{c}$’s are $T_{c}=2|J_{1}-3J_{3}|+6J_{4}$ for the L points, $T_{c}=4|J_{1}+3J_{3}|-12J_{2}-6J_{4}$ for the $\Gamma$ points, and $T_{c}=4J_{2}-6J_{4}$ for the X points.
notation expression $\Gamma(0,0,0)$ X(0,0,$2\pi$) L($\pi,\pi,\pi$) L(others)
$\gamma_{2}({\bm{k}})$ $2\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{z}}{2}\cos\frac{k_{x}}{2}\right)$ $6$ $-2$ 00 00
$\gamma_{4}({\bm{k}})$ $\cos k_{x}+\cos k_{y}+\cos k_{z}$ $3$ $3$ $-3$ $-3$
$\gamma_{1}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\right)$ $4$ 00 $-2$ $2$
$\gamma_{3}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}\zeta^{+}_{{\bm{k}}}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\zeta^{-}_{{\bm{k}}}\right)$ $12$ 00 $6$ $-6$

### v4

# L-point quadrupole order under magnetic field in cubic PrIr2Zn20

@@ Abstract
We study quadrupole orders in heavy-fermion compound PrIr2Zn20 under magnetic fields on the basis of the Landau theory. Assuming $E_{g}$ electric quadrupolar orders in the cubic symmetry with the ordering wavenumber at the L points in the face-centered cubic lattice Brillouin zone as observed experimentally, we construct a Landau free energy and analyze the resulting free energy. We find that the unidentified high-temperature ordered phase under the magnetic field $\bm{H}\parallel[001]$ reported earlier arises from the rotation of the quadrupole moments of f electrons on the Pr site. We also discuss the phase diagram for other magnetic-field directions and possible double-$\bm{q}$ quadrupolar orders in this system.
@@ I Introduction
Strongly correlated electron systems with orbital moments have attracted significant attention in recent years [1, 2]. Various multipolar orders appear in d- and f-electron-based materials and have been intensively studied [3, 4, 5]. Among them, electric quadrupole orders are the most typical and extensively investigated ones [6, 7, 8]. In addition, spin nematic orders have also been studied [9, 10, 11, 12]. The nature of orbital degrees of freedom leads to spatially anisotropic interactions, and thus, the symmetry in orbital space is generally anisotropic. Each type of orbital order has unique properties, and understanding these properties remains a key issue in condensed matter theory and experiment.
@@
Among the 1-2-20 compounds, PrIr2Zn20 is a prototypical material, exhibiting quadrupole ordering below 1 K and undergoing a phase transition to a superconducting state at 0.1 K [31, 32, 33]. The early neutron scattering experiment [34] under an applied magnetic field $\bm{H}\parallel[\bar{1}10]$ suggested an ordering wave number of $\bm{q}=(\pi,\pi,\pi)$, corresponding to the L point at the first Brillouin zone boundary. Here, we have set the lattice constant to unity. The analysis of induced magnetic moments indicated a quadrupole order of the $x^{2}$-$y^{2}$ type for $\bm{H}\parallel[\bar{1}10]$. Recent reinvestigations of the temperature–magnetic field phase diagram have revealed strong anisotropy and highlight the existence of the high-temperature phase [35] reported in earlier studies [31] across a wide range of magnetic field directions from [001] to [111], while it is absent for $\bm{H}\parallel[110]$.
In this paper, we analyze a phenomenological Landau model for PrIr2Zn20, incorporating $E_{g}$ quadrupole moments at the L points. We elucidate the nature of $E_{g}$ quadrupole ordering under magnetic fields and its dependence on field directions. We examine the order parameter suggested by the neutron scattering experiments for $\bm{H}\parallel[\bar{1}10]$, construct phase diagrams under applied magnetic fields, and compare them with the experimental results. The quadrupole orders at the L points reveal unique multiple-$\bm{q}$ orders, which differ from those in the previous studies on multiple-$\bm{q}$ quadrupole orders at the X-point ordering wavenumbers in the face-center-cubic (fcc) lattice [36, 37].
This paper is organized as follows. In Sec. II, we construct the Landau free energy for the $E_{g}$ quadrupole moments at wavenumbers $\bm{q}$$(\pi,\pi,\pi)$ and $(0,0,0)$. In Sec. III, we present the results of the Landau analysis for single-$\bm{q}$ quadrupole orders, outlining the fundamental properties of the system. We then discuss multiple-$\bm{q}$ quadrupole orders and the stability of various ordered phases in Sec. IV. In Sec. V, we compare our results with experimental data for PrIr2Zn20 and related materials. Finally, Sec. VI provides a summary of this paper.
@@ II Model
We begin by introducing the exchange interactions between the $E_{g}$ quadrupole moments in Sec. II.1 and the Landau free energy for the order parameters at the L points in Sec. II.2. Mode-coupling terms related to the order parameter at the L points and other degrees of freedom are discussed in Sec. II.5. Throughout this paper, we do not consider the $T_{xyz}$ octupolar degrees of freedom, since any octupolar orders have not been experimentally reported in PrIr2Zn20 so far. Although the excited CEF states possess various multipole moments such as magnetic dipole, octupole, and $T_{2}$-type quadrupole moments, we will ignore them since they are expected to play a minor role in PrIr2Zn20.
Figure 1: (a) Crystal structure of the diamond structure formed by Pr ions in PrIr2Zn20. AA ($B$) sublattice sites are illustrated in white (red) spheres. The unit cell is taken as two sites along the [111] direction as indicated by shaded oval in green. Exchange interactions $J_{1}\sim J_{4}$ are also indicated. (b) Quadrupole moments $u_{\bm{r}}^{\alpha}$ and $v_{\bm{r}}^{\alpha}$ are schematically shown, where the different color represent the sign of moments.
Figure 2: (a) The first Brillouin zone of the fcc lattice. The four L points are indicated by filled red circles and other symmetric points are indicated by filled black circles. (b)–(d) Fourier transform of the exchange interactions with $J_{1}=1$. (b) $(J_{2},J_{3},J_{4})=(0.0,0.0,0.0)$, (c) $(0.0,0.0,1.0)$, and (d) $(0.5,0.3,0.4)$.
@@ II.1 Exchange interaction
Local moments of f electrons on the Pr sites, which form the diamond structure as shown in Fig. 1(a), tend to order at low temperatures. This is usually caused by Ruderman-Kittel-Kasuya-Yosida interaction due to the presence of the itinerant electrons in heavy-fermion compounds in general. To discuss their ordering, it is reasonable to integrate the itinerant electrons and obtain the effective exchange model for the localized f electron moments. Since the local moments play a dominant role in the symmetry breaking orders in these compounds, we use the localized model as a minimal one to understand the nature of their symmetry breakings. Although it is desirable to obtain the information of the exchange interactions from the excitation spectra by inelastic x-ray or neutron scattering experiments, there is no data available at present. Thus, we use a phenomenological and minimal model which realizes the L point order in this paper.
Let us denote $E_{g}$ quadrupole moments at the position $\bm{r}=(x,y,z)$ in the diamond structure as $\bm{\phi}_{\bm{r}}^{\alpha}=(u_{\bm{r}}^{\alpha},v_{\bm{r}}^{\alpha})$, where $\alpha=A$ or $B$ is the sublattice index. See Fig. 1(b). We take one of the $A$-sublattice site at the origin $\bm{r}=(x,y,z)=(0,0,0)$ and its pair $B$ sublattice site in the unit cell at $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$, where we set the lattice constant $a=1$. $\bm{\phi}_{\bm{r}}^{\alpha}$ forms a two-dimensional vector field as a classical variable. The two components are $u$ $\sim$ $2z^{2}$-$x^{2}$-$y^{2}$ and $v$ $\sim$ $\sqrt{3}$($x^{2}$-$y^{2}$). The Fourier transform in the wavenumber $\bm{k}$} space is defined as
${\bm{\phi}}_{{\bm{k}}}^{\alpha}=\begin{bmatrix}u^{\alpha}_{{\bm{k}}}\\ v^{\alpha}_{{\bm{k}}}\end{bmatrix}=\frac{1}{\sqrt{N}}\sum_{\bm{r}}e^{-i{\bm{k}}\cdot\bm{r}}\begin{bmatrix}u^{\alpha}_{\bm{r}}\\ v^{\alpha}_{\bm{r}}\end{bmatrix}$. (1)
@@
${\mathcal{H}}=\sum_{{\bm{k}}}[u^{A}_{-{\bm{k}}},v^{A}_{-{\bm{k}}},u^{B}_{-{\bm{k}}},v^{B}_{-{\bm{k}}}]\begin{bmatrix}{\mathcal{H}}^{AA}_{{\bm{k}}}&{\mathcal{H}}^{AB}_{{\bm{k}}}\\ {{\mathcal{H}}^{AB}_{{\bm{k}}}}^{\dagger}&{\mathcal{H}}^{AA}_{{\bm{k}}}\end{bmatrix}\begin{bmatrix}u^{A}_{{\bm{k}}}\\ v^{A}_{{\bm{k}}}\\ u^{B}_{{\bm{k}}}\\ v^{B}_{{\bm{k}}}\end{bmatrix}$, (2)
where ${\mathcal{H}}^{\alpha\beta}_{{\bm{k}}}$ is a $2\times 2$ matrix. Such model has been analyzed in incommensurate magnetic orders in the diamond structure [38] and the fcc lattice as a part of the diamond structure [39]. We focus on the orders with the ordering wavenumber at the L points: ${\bm{k}}_{1}\equiv(\pi,\pi,\pi)$, ${\bm{k}}_{2}\equiv(-\pi,\pi,\pi)$, ${\bm{k}}_{3}\equiv(-\pi,-\pi,\pi)$, and ${\bm{k}}_{4}\equiv(\pi,-\pi,\pi)$ as shown in Fig. 2(a). Since the exchange interactions at the L points are isotropic due to the symmetry, we retain such exchange interactions while neglecting anisotropic ones in the following. Although the anisotropic ones are generally present and can influence the ordering vectors and the excitation energy of quadrupole wave, they have no effect on the transition temperature for the order at the L points. Thus, a model without anisotropic interactions can be a good starting model as a minimal one. In this section, we consider exchange interactions up to the fourth neighbor interactions, while we will simplify the model in the later sections.
@@
${\mathcal{H}}^{AA}_{{\bm{k}}}$ $=[J_{2}\gamma_{2}({\bm{k}})+J_{4}\gamma_{4}({\bm{k}})]\hat{1}$, (3)
${\mathcal{H}}^{AB}_{{\bm{k}}}$ $=\frac{1}{2}[J_{1}\gamma_{1}({\bm{k}})+J_{3}\gamma_{3}({\bm{k}})]\hat{1}$. (4)
@@
Here, $\hat{1}$ is a $\times2$ identity matrix. $J_{i}$’s are isotropic coupling constants between the $i$th neighbor sites as indicated in Fig. 1(a). The expressions for $\gamma$’s are listed in Table 1. Note that $\gamma_{1}({\bm{k}})$ and $\gamma_{3}({\bm{k}})$ are complex. From the neutron experimental data, the ordering patterns are consistent with those expected for $J_{1}>0$ [34].
@@
For describing the symmetry breaking at the L-point modulation within our model, it is important to identify relevant interactions. As will be explained later, four special wavenumbers at the $\Gamma$ point and the X points $\{(2\pi,0,0),(0,2\pi,0),(0,0,2\pi)\}$ are important other than the L point. First, among the same sub-lattice interactions, we note that the next-nearest interaction $\gamma_{2}$ vanishes at the four L points, while the fourth neighbor one $\gamma_{4}$ is finite as shown in Table 1. This means that positive $J_{4}$ favors the order at the L points while does not favor those at the $\Gamma$ and X points. For the inter sublattice ones, the nearest-neighbor interaction $\gamma_{1}$ and the third-neighbor one $\gamma_{3}$ are finite, while both vanish at the X points. In this sense, the effects of $J_{1}$ and $J_{3}$ are similar. Diagonalizing the matrix in Eq. (2), one obtains the eigenvalues $J({\bm{k}})$. Figures 2(b)–2(d) show $J({\bm{k}})$ for typical parameter sets with $J_{1}=1$ as a unit of energy. For the nearest-neighbor model with $J_{4}=0$ in Fig. 2(b), $J({\bm{k}})$ is smallest at the $\Gamma$ point, leading to a Néel type antiferoic quadrupole order. When $J_{4}$ is sufficiently large or $J_{2}$ is positive and not too large, the $J({\bm{k}})$ has minima at the L points as shown in Figs. 2(c) and 2(d), and these cases are situations we analyze in this paper. For $J_{1}$-$J_{2}$ model, it is known that the frustration leads to various incommensurate orders in the diamond lattice Heisenberg magnet [40]. Since there is no experimental information about the exchange coupling constants, we try to understand the basic phase diagram for the quadrupole orders in PrIr2Zn20 [31, 32, 35], focusing on the parameter sets that lead to the orders at the L points.
@@
Table 1: exchange interaction coefficients. The values of $\gamma({\bm{k}})$’s at ${\bm{k}}=(0,0,0),(0,0,2\pi)$, ${\bm{k}}_{1}=(\pi,\pi,\pi)$, and the other L points are also shown. Here, we take the unit cell convention of AA: (0,0,0) and $B$: $(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$ and thus, the ${\bm{k}}_{1}$ is inequivalent to the remaining ${\bm{k}}_{2,3,4}$. $\eta_{{\bm{k}}}\equiv e^{i(k_{x}+k_{y}+k_{z})/4}$, $\zeta^{\pm}_{{\bm{k}}}\equiv\pm 2(\cos\frac{k_{x}}{2}+\cos\frac{k_{y}}{2}+\cos\frac{k_{z}}{2})-3$. The mean-field second-order transition temperatures $T_{c}$’s are $T_{c}=2|J_{1}-3J_{3}|+6J_{4}$ for the L points, $T_{c}=4|J_{1}+3J_{3}|-12J_{2}-6J_{4}$ for the $\Gamma$ point, and $T_{c}=4J_{2}-6J_{4}$ for the X points.
[…]
$\gamma_{2}({\bm{k}})$ $2\left(\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{z}}{2}\cos\frac{k_{x}}{2}\right)$ 66 2-2 0 0
$\gamma_{4}({\bm{k}})$ $\cos k_{x}+\cos k_{y}+\cos k_{z}$ 33 33 3-3 3-3
$\gamma_{1}({\bm{k}})$ $\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\right) 44$ 0 2-2 22
$\gamma_{3}({\bm{k}})$ $4\eta_{{\bm{k}}}\left(\cos\frac{k_{x}}{4}\cos\frac{k_{y}}{4}\cos\frac{k_{z}}{4}\zeta^{+}_{{\bm{k}}}+i\sin\frac{k_{x}}{4}\sin\frac{k_{y}}{4}\sin\frac{k_{z}}{4}\zeta^{-}_{{\bm{k}}}\right)$ 1212 0 66 6-6